OpenStudy (anonymous):

***Answer for a medal*** Planet A takes 1 year to go around its star at an average of 1 A.U. distance. Planet B is 4 A.U. from the star. Calculate how long Planet B takes to orbit. ______________ years 512 64 23 8

OpenStudy (anonymous):

can i get some help please :'(

OpenStudy (anonymous):

I guess the answer is 8. by kepler's third law, \[T ^{2} \alpha A^{3}\] T=time in years A=Astronomical Unit. as A=4 therfore \[A ^{3}=64\] and \[T ^{2}=64 \] hence T=8 But it is not the correct answer you should try the formula\[T^{2}=\left( \frac{ 4\pi ^{2} }{ GM } \right)A ^{3}\] where M=mass of the Sun=\[2\times10^{30}kg\] and G is gravitational constant=\[6.67\times10^{-11}Nm ^{2}/kg ^{2}\] 1AU=\[1.5\times10^{8}km\] I tried it but there must some calculation error, Best of luck for you.

OpenStudy (anonymous):

\[T^2 \to a^3\] - kepler's law of planetary motion. so around the same star, let orbital parameters for planet A be T1 , a1 and for Planet B, T2, a2 hence, \[T2^2/T1^2 = A2^3/A1^3\] T1 = 1, A1 = 1, A2 = 4 hence, \[T2 = \sqrt{64} = 8\] in years